OpenStudy (anonymous):

okay so in math today we learned how to write equations in standard for and the list the a , b , c value can you see if i got this right : 2x^2+28-5x=0 you just rewrite it and the a=2 b=28 and c= -5 can yall tell me if i got it right !!

OpenStudy (anonymous):

not quite.. your a is right... b is the coefficient of the x term, and c is the constant coeffiecient

OpenStudy (anonymous):

the equation is : ax^2 + bx + c = 0

OpenStudy (anonymous):

in standard form...

OpenStudy (anonymous):

wait im still confuzed

OpenStudy (australopithecus):

a, and b are the coefficients in front of the variable in this case x is the variable: ax^2 + bx + c = 0 thus 3x^(2) + x + 2 = 0 we have that a = 3 b = 1 c = 2

OpenStudy (australopithecus):

Another example -2x^(2) + 2 = 0 a = -2 b = 0 c = 2 Hint: Look at your answer again

OpenStudy (australopithecus):

One last example - x^(2) - 4x = 0 a = - 1 b = - 4 c = 0

OpenStudy (australopithecus):

c is just a number with no variable multiplied by it